### v1

# Efficient Multilingual Search Relevance Modeling in E-Commerce via LLM Mixture-of-Experts

Abstract.
In e-commerce platforms, search relevance directly influences both user experience and merchant revenue. In multi-country deployments, diverse linguistic, cultural, and product catalog contexts introduce significant distribution shifts, posing substantial challenges to relevance modeling. Existing approaches typically enhance the reasoning or multilingual abilities of a single monolithic model, yet they remain limited by data diversity, coverage gaps, and high inference costs in heterogeneous environments. Our empirical analysis reveals that different LLM base models exhibit complementary strengths across languages and regions, motivating an expert-based architecture. We propose a scalable LLM-based Mixture-of-Experts (MoE) framework that dynamically routes queries to specialized experts and fuses their embeddings through concatenation. Among rule-based, pseudo-label–based, and fully end-to-end strategies, end-to-end hard routing with concatenation offers the best balance of effectiveness and efficiency. To mitigate inference overhead, we further develop an engineering-optimized offline batch pipeline with resource-efficient scheduling, which hides memory latency, improves GPU utilization, and reduces GPU-hour consumption by up to 35% compared with synchronous execution. On datasets spanning six Southeast Asian markets, our MoE improves AUC by 0.72 percentage points over a dense baseline with the same active parameters. Meanwhile, the optimized pipeline achieves 27.6 queries per second (QPS), a 9% throughput improvement. These results demonstrate superior multilingual relevance and efficiency, delivering strong cost-effectiveness for real-world e-commerce search systems.
E-commerce Retrieval, Search Relevance, Large language models
copyright: acmlicenseddoi: XXXXXXX.XXXXXXXconference: The ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesjournalyear: 2026isbn: 978-1-4503-XXXX-X/2018/06submissionid: 210ccs: Applied computing Online shoppingccs: Computing methodologies Natural language generationccs: Information systems Query intent
Figure 1. Radar chart showing AUC of three base models across multiple country–language tasks. :
1. Introduction
In recent years, the rapid growth of e-commerce platforms across multiple countries has made search relevance in multilingual and multicultural environments a critical factor affecting both user experience and merchant revenue. Traditional approaches typically employ BERT-like models to learn contextualized representations of queries and product descriptions, determining relevance via semantic similarity. With the advent of large language models (LLMs), their extensive knowledge and strong language understanding have opened a new paradigm for relevance estimation.
Most existing LLM-based methods focus on enhancing the capabilities of a single model, such as improving reasoning through fine-tuning (10.1145/3701716.3715246; zhao2025explainable), incorporating multimodal signals (Zheng_2023; zhu2023querylifequeryawarelanguageimage) like product images or user behavior, or performing continued pre-training (dou2025sailor2sailingsoutheastasia) to strengthen multilingual understanding. While these approaches can achieve notable gains, they often demand substantial training and inference resources, and collecting sufficiently diverse and representative data across multiple real-world markets remains challenging. Such single-model enhancements can be directly integrated into our proposed framework as independent experts. Our empirical analysis (Figure 1) shows that even with identical fine-tuning data/objectives, different LLMs (Gemma2-9B (gemma_2024), Qwen2.5-14B (qwen2.5), Sailor2-20B (dou2025sailor2sailingsoutheastasia)) vary in performance across languages/regions. For example, Sailor2-20B (larger scale, Southeast Asian languages pre-training) does not outperform others in some languages, while Qwen2.5-14B excels in Indonesian–English and Singapore–Vietnamese. This suggests that differences in pre-training corpora, vocabulary design, and language coverage lead each model to internalize distinct regional language capabilities, which remain complementary even under identical fine-tuning.
Building on these observations, we propose a lightweight LLM-based Mixture-of-Experts (MoE) framework that integrates fine-tuned LLMs as independent experts without additional fine-tuning. The routing module dynamically assigns each input to the most suitable experts, and their representations are fused at the embedding level, leveraging both the improvements within each expert and the complementary strengths across models. This design exploits the advantages of multiple base models in multilingual and multi-regional relevance tasks, achieving a balanced trade-off between effectiveness and efficiency. We explore various routing strategies, including rule-based, pseudo-label, and end-to-end (soft/hard) approaches, and find that end-to-end hard routing delivers the best overall balance. Experiments on real-world multi-country e-commerce datasets show that our method improves AUC by 0.6 percentage points and throughput by approximately 120% over a strong dense baseline, while maintaining low inference latency.
2. Methodology
To leverage the complementary strengths of LLMs in multilingual e-commerce search while controlling inference cost, we propose a sparsely-activated Mixture-of-Experts (MoE) framework—where each LLM acts as an independent expert. This framework jointly optimizes expert selection and representation fusion to balance high relevance performance with predictable computational overhead.
Figure 2. Overview of the proposed sparsely-activated Mixture-of-Experts (MoE) framework for multilingual e-commerce search. The routing module dynamically selects top-k LLM experts based on query, nation, and item title; their hidden representations are projected and concatenated in the fusion module for relevance prediction. The bottom part shows the architectures of the routing and fusion modules, including soft and hard routing strategies.
2.1. LLM-based Mixture-of-Experts Architecture
Our framework integrates multiple finetuned LLMs as experts. Given a query–item pair, a routing module first selects a subset of experts, whose contextual embeddings are then fused for final relevance prediction. The routing module balances prediction performance and computational efficiency. We consider four strategies:
Rule-based routing assigns each query to a single expert based on handcrafted criteria (e.g., language). It is non-parametric, incurs no additional cost, but is rigid and limited to top-1 selection.
Pseudo-label routing uses a two-stage pipeline: first, LLMs infer pseudo-labels for the best-performing expert per sample; then a lightweight router is trained to predict these labels, followed by training the fusion module to align models.
End-to-end soft routing outputs a probability distribution over all $N$ experts and applies an entropy minimization regularizer to encourage sparsity: $L_{\text{ent}}=\sum_{i=1}^{N}p_{i}\log p_{i}$, where $p_{i}$ is the routing probability for expert $i$. This encourages sharper expert selection and model specialization. During inference, only experts exceeding a threshold are activated. Soft routing may suffer from threshold sensitivity and train–inference mismatch.
End-to-end hard routing directly selects the top-$k$ experts (e.g., $k=2$) and incorporates a load balancing loss to prevent routing collapse. Differences in pre-training corpora and architectures produce heterogeneous hidden representations across experts. Early in training, the fusion module may overfit to the easiest expert, causing vanishing gradients for the router. To mitigate this, we adopt a load balancing loss (fedus2022switchtransformersscalingtrillion):
(1) $L_{\text{LB}}=N\cdot\sum_{i=1}^{N}p_{i}\cdot\bar{p}_{i},$
where $p_{i}$ is the fraction of samples routed to expert $i$, and $\bar{p}_{i}=1/N$ is the ideal uniform usage. This stabilizes training, ensures diverse expert utilization, and maintains consistency between training and inference. Hard routing also provides predictable latency and avoids threshold tuning, which we adopt as the default.
Table 1. Overall and per-market AUC (%) and QPS for all baselines and our proposed MoE. Bold indicates the best, underline the second-best.
AUC ($\uparrow$) QPS ($\uparrow$)
Model Overall ID MY PH SG TH VN
Qwen2.5-14B 91.40 86.08 92.44 88.99 93.62 93.82 93.41 28.34
Gemma2-9B 91.40 86.35 93.71 89.49 93.09 94.00 92.94 36.76
Qwen+Gemma Fusion 91.88 86.76 93.47 89.68 93.61 94.27 93.74 16.01
Sailor2-20B 91.82 87.63 92.40 89.24 93.14 94.65 93.64 21.90
Qwen2.5-32B 91.71 86.43 92.57 89.58 93.83 94.08 93.55 12.55
Full Fusion (Concat) 92.41 87.89 93.70 90.25 93.75 94.94 94.13 6.94
Proposed MoE 92.49 88.33 93.56 90.45 93.62 95.00 94.10 13.72
For each selected expert $i$, we extract its final transformer layer’s last-token hidden state $\mathbf{h}_{i}\in\mathbb{R}^{d_{i}}$, then project it to a shared space:
(2) $\mathbf{h}^{\prime}_{i}=W_{i}\mathbf{h}_{i}+\mathbf{b}_{i},\quad\mathbf{h}^{\prime}_{i}\in\mathbb{R}^{d}.$
The projected embeddings are then concatenated:
(3) $\mathbf{h}_{\text{concat}}=[\mathbf{h}^{\prime}_{i_{1}}\parallel\dots\parallel\mathbf{h}^{\prime}_{i_{k}}]\in\mathbb{R}^{kd}.$
Concatenation preserves non-redundant knowledge (e.g., syntax, cultural context) from multiple experts, unlike weighted fusion which may prematurely compress features. The fused vector is passed through a lightweight classifier:
(4) $\hat{y}=\sigma(W_{c}\mathbf{h}_{\text{concat}}+\mathbf{b}_{c}).$
The training loss combines cross-entropy and load balancing:
(5) $L=L_{\text{CE}}+\lambda L_{\text{LB}},$
where $\lambda$ balances task performance and expert diversity.
2.2. Offline Batch Inference with Resource-Efficient Scheduling
Integrating multiple large LLMs as independent experts introduces substantial computational overhead during inference, particularly in real-world e-commerce environments with high query volumes and strict latency requirements. To address this, we design an offline batch inference pipeline that leverages multi-stream asynchronous execution and resource-efficient scheduling, ensuring high throughput while preserving relevance accuracy.
Design Principles. The pipeline is built on three principles: (1) sparsity-aware computation, activating only a top-$k$ subset of experts per query; (2) parallel execution, overlapping computation and memory operations across experts to exploit GPU-level concurrency; and (3) scalable scheduling, dynamically allocating resources to balance load among heterogeneous experts.
Three-Stage Workflow. Inference proceeds in three decoupled stages. (1) Bulk Routing Stage: The router processes queries in large batches and selects the top-$k$ experts for each query based on query text, region, and item metadata. By precomputing or caching routing decisions where applicable, this stage introduces minimal overhead and enables sparse activation of experts, which reduces memory and compute requirements. (2) Expert-Specific Batch Inference Stage: Queries assigned to the same expert are grouped into batches for parallel execution. Each expert independently performs forward propagation on its batch, and multiple experts can run concurrently on separate devices or asynchronously on the same device. This stage exploits GPU parallelism and allows dynamic resource allocation, where faster experts can assist slower ones, improving cluster-level utilization and overall throughput. (3) Late-Stage Fusion: After expert inference, the hidden states are aggregated and fused according to the selected fusion strategy (e.g., projection and concatenation). This stage involves only lightweight linear transformations and a small classifier, and can be efficiently executed on lower-cost devices. The design ensures that heterogeneous representations from different LLMs are preserved and combined without loss of complementary knowledge.
3. Experiments
3.1. Dataset and Evaluation Metrics
We construct a large-scale multilingual e-commerce relevance dataset, collected from real-world Alibaba Lazada search logs and annotated by professional evaluators, covering six Southeast Asian markets: Indonesia (ID), Malaysia (MY), Philippines (PH), Singapore (SG), Thailand (TH), and Vietnam (VN). Each query–item pair is labeled as Yes (relevant) or No (irrelevant). The dataset comprises 7.1M training samples, 0.7M validation samples, and 46k test samples. We use the Area Under the ROC Curve (AUC) as the primary effectiveness metric and Queries Per Second (QPS) as the efficiency metric. QPS is measured on an NVIDIA H20 GPU and averaged over 1,000 consecutive batches to ensure stable evaluation.
3.2. Experimental Settings
Base models—Qwen2.5-14B, Gemma2-9B, Sailor2-20B, and Qwen2.5-32B—are fine-tuned on the same training set using RSLora(kalajdzievski2023rankstabilizationscalingfactor) with rank $r=256$ and scaling factor $\alpha=128$. Training is conducted for a single epoch with batch size 4 and learning rate $5\times 10^{-6}$, following standard practice for large-scale relevance fine-tuning. The MoE framework, including the router, projection layers, and fusion classifier, is trained with batch size 128, learning rate $1\times 10^{-4}$, and $\lambda=0.01$ in Eq. 5, while all base LLMs remain frozen.
We include Qwen2.5-32B as a strong dense baseline, whose total parameter count (32B) is comparable to the maximum active parameters of our MoE framework ( 34B for two experts). This enables a fair comparison between a scaled-up homogeneous model and our sparsely-activated heterogeneous MoE. Other baselines include smaller individual models and a full fusion (Concat) variant that concatenates all three experts’ embeddings without routing.
Table 2. Ablation studies on fusion and routing strategies. Reported metrics are overall AUC (%) and QPS.
Method AUC (%) $\uparrow$ QPS $\uparrow$
Fusion strategies (Hard routing)
Weighted fusion 92.27 13.72
Concatenation 92.49 13.72
Routing strategies (Concat fusion)
Rule-based 91.94 24.44
Pseudo-label 91.57 34.47
Soft routing 92.09 16.01
Hard routing 92.49 13.72
Serial Hard routing 92.49 6.56
3.3. Main Results
Table 1 reports overall and per-market AUCs and QPS. Our MoE achieves the highest overall AUC (92.49%), surpassing all single models and the full fusion variant—proving its strong multilingual relevance modeling. Across markets, it integrates expert strengths: best performance in ID/PH/TH, and competitive scores in MY/SG/VN. Notably, it outperforms the full fusion model in several regions despite activating fewer experts, suggesting that full fusion may introduce noise and limit effective use of each base model. In terms of efficiency, the MoE achieves a QPS of 13.72, exceeding the dense Qwen2.5-32B (12.55) and substantially outperforming full fusion (6.94), achieving a favorable relevance-throughput trade-off. We also observe that Sailor2-20B, obtained by extending and further pretraining Qwen2.5-14B, improves substantially over its base model. Nevertheless, even a simple fusion of Qwen and Gemma yields higher overall performance than Sailor2-20B, with much lower pretraining cost. This indicates that combining complementary experts through lightweight fusion can be more effective and efficient than scaling a single model via continued pretraining.
3.4. Ablation Studies
To assess the contribution of individual components, we conduct ablation studies on fusion strategies and routing mechanisms. All variants are trained and evaluated under identical settings. Table 2 shows concatenation-based fusion outperforms weighted fusion by 0.22 AUC, as it preserves complementary features without premature compression. Among routing strategies, end-to-end hard routing achieves the highest effectiveness, whereas pseudo-label routing attains the highest QPS but performs poorly because most samples collapse to Gemma during training due to difficulty in learning accurate pseudo labels. Soft routing suffers from a train–inference mismatch, leading to suboptimal performance. We also include a Serial Hard routing variant, where selected experts are processed sequentially rather than in parallel. While it achieves the same AUC as hard routing, the QPS drops significantly (from 13.72 to 6.56), demonstrating that our optimized parallel inference pipeline nearly doubles throughput while maintaining identical model outputs. This highlights the effectiveness of the resource-efficient scheduling and multi-stream batch execution in practical deployment.
3.5. Online Deployment
In industrial search advertising systems, upon receiving a user query, a Top-K subset of ads is selected from a vast inventory, following a standard multi-stage pipeline: index → retrieval → prerank → rank → auction → rerank. At the retrieval stage, query–item relevance scores serve dual purposes: filtering candidates for downstream stages and informing dynamic reserve prices in the auction mechanism (feng2018deep; Li_2025), which balance user experience, platform revenue, and advertiser ROI. Accurate relevance estimation is therefore critical for both result quality and monetization efficacy.
However, industrial-scale retrieval involves tens to hundreds of thousands of candidate items per query. Real-time inference with large language models (LLMs) at this scale is computationally prohibitive due to latency constraints. To leverage LLM semantics without losing efficiency, we use offline knowledge distillation: LLMs generate high-quality relevance labels on a billion-scale query-item corpus to train a compact student model. We select ColBERT (khattab2020colbertefficienteffectivepassage) as the student model due to its strong representational capacity and efficiency in late-interaction architectures. The distilled ColBERT model is deployed online, enabling fine-grained, low-latency relevance scoring that meets the demands of real-time retrieval.
It should be noted that the translation of improved AUC into actual gains in ad revenue is influenced by multiple confounding factors, such as dynamic reserve price parameters and shifts in the online bidding environment. To ensure rigorous interpretation and avoid potential misattribution, we refrain from reporting online experimental results at this stage.
4. Conclusion
We propose an LLM-based Mixture-of-Experts (MoE) framework for multilingual e-commerce search relevance, leveraging LLMs’ cross-lingual/regional complementary strengths via dynamic routing and embedding fusion. Experiments on six Southeast Asian markets show our MoE boosts AUC by 0.72 percentage points over dense baselines. Our optimized offline pipeline reduces GPU-hours by 35% and raises QPS throughput by 9%. End-to-end hard routing with concatenation best balances effectiveness and efficiency. It delivers practical value for real-world deployment; future work will explore multimodal integration and automated expert selection.